### v1

# Generative AI for Intent-Driven Network Management in 6G: A Case Study on Hierarchical Learning Approach

Abstract
With the emergence of 6G, mobile networks are becoming increasingly heterogeneous and dynamic, necessitating advanced automation for efficient management. Intent-Driven Networks (IDNs) address this by translating high-level intents into optimization policies. Large Language Models (LLMs) can enhance this process by understanding complex human instructions to enable adaptive, intelligent automation. Given the rapid advancements in Generative AI (GenAI), a comprehensive survey of LLM-based IDN architectures in disaggregated Radio Access Network (RAN) environments is both timely and critical. This article provides such a survey, along with a case study on a hierarchical learning-enabled IDN architecture that integrates GenAI across three key stages: intent processing, intent validation, and intent execution. Unlike most existing approaches that apply GenAI in the form of LLMs for intent processing only, we propose a hierarchical framework that introduces GenAI across all three stages of IDN. To demonstrate the effectiveness of the proposed IDN management architecture, we present a case study based on the latest GenAI architecture named Mamba. The case study shows how the proposed GenAI-driven architecture enhances network performance through intelligent automation, surpassing the performance of the conventional IDN architectures.
Index Terms:
Intent processing, Intent validation, Intent execution, Generative AI, Hierarchical learning
I Introduction
Sixth-Generation (6G) networks are anticipated to support a diverse set of user requirements and have more complex deployments [1]. Traditional network management methods, which are based on manual configurations and human expertise, will face challenges when optimizing and operating such complex networks. Furthermore, manual configurations are costly, prone to errors and not scalable [2]. Intent-Driven Network (IDN) management is emerging as a solution to these challenges [1]. An intent defines desired outcomes by specifying target metrics without detailing implementation steps to ensure both human comprehension and machine interpretability [4]. Advances in Natural Language Processing (NLP) help operators express these intents more effectively. This reduces manual work, lowers errors, and speeds up configuration changes and service deployments [4, 5].
Disaggregated Radio Access Networks (RANs), like Open RAN with its layered structure, are well-suited for IDN management. High-level controllers can interpret operator goals and check if the intent is achievable based on real-time conditions. Low-level controllers or functions then create the best network policies to meet these goals. This approach improves flexibility, scalability, and decision-making efficiency.
On the other hand, Generative Artificial Intelligence (GenAI) has led to the rapid advancement of Large Language Models (LLMs) and revolutionized the field of NLP. As they have an outstanding ability to understand complex human instructions in natural language, IDN management is leaning towards LLMs instead of using fixed Service Level Agreement (SLA)-based approaches. For example, Dzeparoska et al. have introduced a pipeline that utilizes an LLM to process intents into structured and policy-based abstractions, which are next linked with Application Programming Interfaces (APIs) for execution [6]. Also, in our previous work we have introduced an IDN management framework that uses LLM few-shot learning for intent processing from a human network operator [4]. While these works use LLMs for intent processing, it can be highly beneficial to introduce GenAI not only as an intent processor via LLM but also we can take IDN management to the next level by introducing GenAI solutions in the other two stages namely intent validation and intent execution.
Figure 1: Three-step methodology for intent-driven network management.
With the inspiration for introducing efficient GenAI-based IDN management schemes via hierarchical disaggregated RAN environments, this work first reviews the state-of-the-art approaches that use GenAI for IDN management. Next, we introduce a three-step methodology for IDN management, encompassing intent processing, validation, and execution (see Fig. 1 for details). After that, we present how these three key steps can be implemented using a hierarchical RAN architecture with the help of multiple GenAI algorithms. Lastly, we present a case study that shows the implementation of these steps. In particular, we process intents in a memory-efficient manner where an LLM is fine-tuned with a custom dataset using Quantized Low Rank Adaptation (QLoRA) [7]. For the intent validation, a transformer-based time series predictor has been used in the case study which predicts crucial network parameters in future time slots and based on the predictions validates intents. Lastly, for intent execution, we generate network policies using Mamba, a GenAI architecture based on state space models for its demonstrated effectiveness compared to the transformer architecture in terms of memory and computation efficiency [8].
The core contributions of this paper are as follows: it presents a precise, yet comprehensive survey of LLM-based IDN architectures, highlighting current limitations and identifying opportunities for integrating broader GenAI techniques across the stages of intent processing, validation, and execution. Furthermore, it introduces a three-step methodology for IDN management that encompasses these stages, integrated within a hierarchical RAN architecture. To validate the proposed framework, a case study demonstrating the utility of the proposed framework using decision Mamba has been presented for the first time in the literature.
II GenAI-based Intent-driven Network Management Schemes
Intent-based networking connects end-user service requests with network operations. Nijah et al. states that NLP-based methods offer the most flexibility to enable intent expression in natural language [2]. LLMs have revolutionized NLP and they surpass traditional methods in flexibility and accuracy. Their ability to perform zero-shot, few-shot, and fine-tuned learning makes them state-of-the-art for tasks like intent recognition and processing. Existing research on IDN management schemes vastly uses LLMs to understand intents from the operators. While a few GenAI paradigms explore techniques like diffusion models and neuro-symbolic reasoning, the majority of current research in IDN management focuses on LLMs. Therefore, in this section, we review the LLM-centric approaches for IDN management in the telecommunication domain. Subsequently, we explore the potential application of LLM techniques or GenAI algorithms which to date, have not been widely explored in the literature. We also summarize the LLM-centric IDN management approaches, their main features, challenges, and applications in Table I.
TABLE I: LLM-Based IDN Management Approaches: Features, Challenges, and Applications
Category Main Features Challenges Applications
LLM-assisted intent processing and management for Fifth-Generation (5G) core networks [9, 10] Supports multiple intent transmission to the model. Enhances network orchestration with semantic routing. Prevents LLM hallucination. Includes a RAG module. Reliance on LLMs which are not inherently designed for network control and automation. Distinguishing between similar or overlapping intents. Conversion of operator-defined intents into slice deployment policies. Mapping intents to edge-based service deployments.
LLM-Driven multi-agent and negotiation-based intent management for 6G networks [11, 5] Employs MARL and LLM-driven negotiation. Resolves conflicting intents in multi-stakeholder settings. Cross-domain conflict resolution. Fairness in resource allocation. Scalability concerns. 6G network slicing. Dynamic spectrum allocation. Autonomous resource management.
End-to-end AI-based intent-driven network automation [1, 2, 4] Converts high-level intents into actionable network policies. Verifies if an intent is feasible given the network’s current or predicted state. End-to-end automation using AI. Computational complexity of processing and validating multiple intents in real-time increases exponentially. Zero-touch network configuration. Predictive maintenance and fault recovery.
Generative AI and multimodal intent-based network management [12] Uses multimodal generative AI for intent translation. LLMs are fine-tuned with few-shot learning to adapt to networking use cases. High dependency on predefined templates. Expansion to include multimodal inputs (e.g., images, topology files). Automation of the deployment and management of network slices.
II-A State-of-the-art LLM-Based Approaches for IDN Management
II-A1 LLM-Assisted Intent Processing and Management for 5G Core Networks
Two recent studies explore LLM-assisted approaches on intent-based management in 5G core networks. Manias et al. in their work [9] explore LLM-driven intent extraction as a key enabler of next-generation zero-touch network service management. A customized LLM model is designed to interpret and translate user intents into actionable network policies. By utilizing LLMs, the framework reduces human intervention while maintaining adaptability across diverse network management tasks, such as deployment, modification, performance assurance, and feasibility checking.
Semantic routing is introduced to refine LLM-assisted intent-based networking in [10]. Traditional LLM-driven approaches face issues such as hallucination, scalability limitations, and accuracy degradation when handling complex network intents. To address these challenges, the proposed framework integrates a semantic router, which ensures deterministic decision-making by routing extracted intents through predefined pathways rather than relying solely on LLM-generated responses. This hybrid architecture improves reliability, reduces latency, and enhances accuracy in 5G core management and orchestration.
II-A2 LLM-Driven Multi-Agent and Negotiation-Based Intent Management
The framework proposed in [11] is a collaborative, multi-agent system for managing shared network resources in 6G. This system deploys LLM-based agents to represent different business entities. Each of these entities negotiates service-level objectives such as throughput, cost efficiency, and energy savings. The framework acts as a central mediator to resolve conflicts by utilizing LLMs alongside optimization techniques and real-time network observability.
Another study [5] introduces a comprehensive LLM-driven intent life cycle management system designed to handle all stages of intent processing, from decomposition and translation to negotiation, activation, and assurance. By shifting away from rigid JSON/YAML intent definitions, the system proposed in [5] enables natural language-driven network orchestration. Users can specify operational goals without requiring deep technical expertise.
II-A3 End-to-End AI-Based Intent-Driven Network Automation
Recent studies introduce end-to-end AI frameworks that integrate LLMs, algorithms like Deep Reinforcement Learning (DRL), Multi-Agent RL (MARL), and Hierarchical RL (HRL) to realize fully automated network management.
While processing intents is crucial, ensuring their feasibility and impact is equally important [4]. A recent study introduces a transformer-based time series predictor to validate intents before execution. This predictive validation module analyzes historical network data and forecasts traffic patterns to ensure that requested optimizations (e.g., increasing energy efficiency, improving throughput) do not negatively affect service quality. Once an intent is validated, an HRL framework selects and triggers appropriate network optimization applications such as beamforming, traffic steering, and power control. An attention-based HRL model filters out suboptimal actions to reduce computational overhead while maximizing efficiency.
Another breakthrough in end-to-end AI-driven network automation is the integration of multi-agent learning frameworks [1], where AI-driven network agents negotiate and resolve conflicting intents in real-time. This is particularly essential in multi-tenant 6G networks, where various stakeholders (e.g., operators, service providers, enterprise users) have competing resource demands. The study conducted in [1] proposes a multi-domain orchestration model, where LLM-driven agents representing RAN, core, and transport domains collaborate to optimize resource allocation and maintain Quality-of-Service (QoS).
II-A4 Generative AI and Multimodal Intent-Based Network Management
A multimodal intent recognition framework presented in [12] utilizes LLMs with GenAI to process diverse inputs like text, images, and deployment descriptors to enable network operators to specify intents through text prompts, graphical designs, or configuration files. AI dynamically refines intents to ensure accuracy while aligning with service-level objectives.
A key application of this approach is network-as-a-service orchestration, where LLM-powered agents automate network slice management. The framework enables zero-touch deployment to map business requirements to network slice templates.
II-B Potential GenAI Paradigms for Enhancing IDN Management
The key limitations of the LLM-based IDN management schemes that we discussed in this section can be described as follows: LLM hallucination and interpretability challenges, which can result in inaccurate or overly generalized outputs; absence of a real-time feedback mechanism, making it difficult to correct errors in network policies after deployment; and deployment challenges, particularly due to memory constraints and high computational requirements. To address these issues, we recommend some possible solutions from the GenAI perspective.
$\bullet$ Retrieval Augmented Generation (RAG) and separate intent refinement/ validation scheme: We can use network-specific databases and retrieval models to ensure LLMs only generate responses based on verified knowledge to reduce hallucination. Furthermore, a multi-step intent refinement process can be introduced where a second model (e.g., a smaller LLM or rule-based checker) cross-checks intent translations before execution.
$\bullet$ Reinforcement Learning with Human Feedback (RLHF): Since re-training is not always an option due to computational resource scarcity, RLHF can be a powerful technique for improving LLM-based IDN management to address the issue of real-time adaptation and feedback loops. We can define a reward function based on network Key Performance Indicators (KPIs) (e.g., latency, throughput, energy efficiency). Then, a feedback loop can be implemented where operators and simulated environments score the LLM ’s intent outputs. Lastly, RL algorithms can be used to refine LLM-generated responses.
$\bullet$ LLM quantization and compression: Computation overhead caused by LLM deployments can be handled by implementing 4-bit quantization techniques (QLoRA, Generative Pre-trained Transformer-Quantization (GPTQ)) to reduce LLM model size while maintaining high accuracy.
III Three-step Methodology for Intent-driven Network Management
In this section of the paper, we present a three-step framework to introduce IDN management in modern-day communication systems. In particular, we propose a three-fold strategy consisting of intent processing, intent validation, and intent execution via network optimization policy generation.
Figure 2: Integration of the proposed IDN management framework with hierarchical RAN architecture.
$\bullet$ Intent processing: Intent processing is the phase where high-level, human-expressed network intents (e.g., “Minimize latency for video traffic”) are interpreted, extracted, and translated into structured, machine-executable commands using LLMs, multimodal AI, or rule-based systems. It is the first step of the IDN management and is highly important because if intent interpretation is not done correctly, it can lead to performance degradation.
$\bullet$ Intent validation: Intent validation is the process of assessing whether the current network conditions and capacity can accommodate a given intent or a set of intents. This involves verifying whether the available resources, traffic conditions, and operational constraints allow for the successful execution of the desired network optimization actions. Intent validation ensures that a processed intent is feasible, conflict-free, and optimal before execution.
$\bullet$ Intent execution: Intent execution is the final stage of IDN management, where validated intents are transformed into real-time network actions. This phase involves invoking appropriate network optimization policies to dynamically configure resources, and ensure that the desired service-level objectives are met efficiently. Intent execution can be realized through network function orchestration, AI-driven decision-making, or Reinforcement Learning (RL)-based adaptation. There are multiple ways to execute intents depending on the network architecture and management framework. The system can trigger predefined network applications such as traffic steering, beamforming, power control, and slicing management for intent execution. Also, LLMs can generate and refine network policies dynamically. Another approach for multi-operator environments can be to utilize RL-based agents that can negotiate optimal execution strategies to ensure efficient network optimization.
To represent this three-step methodology, we presented Fig.1 in the Introduction, which illustrates the step-by-step process of intent-driven network management. The process begins with intent processing, where network intents can be expressed in natural language, structured formats (JSON, XML), or multi-modal data such as topology diagrams and configuration files. To process the intents, different methodologies can be applied, including LLMs for natural language understanding, rule-based approaches using predefined logic, graph-based methods for structured representation of dependencies, or finite state machines to model intent workflows as a sequence of state transitions.
Once an intent is processed, it moves to intent validation which can be carried out using rule-based validation (checking against static policies), constraint-based validation (ensuring resource availability and compliance), predictive validation (using AI/ML models to anticipate network conditions), or RLHF to refine validation mechanisms over time.
After successful validation, the intent execution phase applies the validated intent to the network infrastructure. The execution can be performed using API-based deployment for direct implementation, Open RAN network applications for network-specific functionalities, or negotiation-based execution via multiple LLM agents to resolve potential conflicts between different intents before deployment.
Figure 3: End-to-end GenAI framework for intent-driven network management.
III-A Hierarchical RAN Integration of the Proposed IDN Management Architecture
Modern hierarchical RAN architectures, such as Open RAN, use a multi-layered and modular design to support flexibility, scalability, and interoperability in next-generation wireless networks. These systems break down the traditional monolithic RAN into software-driven components that enable automation and AI-based optimization. In this work, we adopt a two-level architecture composed of a strategic controller and a tactical controller. The strategic controller manages long-term objectives like service quality, energy efficiency, and traffic patterns, while the tactical controller focuses on real-time operations including handovers, resource allocation, and traffic steering. This layered approach helps implement intent processing, validation, and execution by supporting structured decisions across different time scales and abstraction levels, ensuring accurate interpretation and efficient execution of network intents.
We provide an example of implementing the three-step IDN management workflow using a custom hierarchical RAN architecture in Fig. 2. A human operator can observe the current network status or query for any relevant network information before providing an intent. The LLM module in the picture is given as an example. The operator may use logs, or another query-based database to gather information instead of using an LLM. After that, the operator can provide an intent that is received by the user interface application in the top-level Intelligent RAN Controller (IRC) named strategic controller. It then passes through the intent validation application. This application can be designed to perform intent validation via rule-based static methodologies or GenAI-based predictive analysis. Lastly, at the bottom, there is a tactical controller responsible for intent execution. In Fig. 2, we provide an example of intent execution via network applications in the tactical controller. Based on the intent, multiple network applications such as traffic steering, power allocation, and cell sleeping can be initiated and orchestrated to fulfill an intent.
IV A Case Study with Hierarchical Learning for IDN Management
Hierarchical application of GenAI algorithms can optimize complex decision-making by structuring learning at multiple levels. For instance, LLMs can interpret human intents, transformer-based predictors can validate network conditions, and decision transformers can execute intents. At the highest level, an LLM extracts key metrics and constraints from operator objectives. These insights guide a mid-tier forecasting model that predicts network behavior using historical and real-time data. At the lowest level, control algorithms generate policies for intent execution. This layered hierarchical approach improves efficiency, reduces computational overhead, and enables adaptive, real-time decisions
Fig. 3 presents an end-to-end GenAI framework for IDN management, which forms the foundation of our case study. Data collection follows a hierarchical reinforcement learning approach with decision-making split across two levels. The higher-level controller defines goals based on operator intents, such as increasing throughput or reducing power usage. The lower-level controller selects appropriate network applications depending on real-time conditions like traffic load, energy consumption, and packet loss. Five applications are used in this study: traffic steering, cell sleeping, beamforming, power allocation, and energy-efficient handover management. Each of them is developed using DRL. Traffic steering application distributes traffic intelligently across Radio Access Technologys (RATs) to maintain QoS, optimizing both delay and throughput. Cell sleeping reduces power consumption by deactivating underutilized base stations based on traffic load and queue length. Beamforming calculates optimal steering angles and selects matching array vectors using user location and the power allocation application improves total throughput through smart resource distribution. Lastly, the handover management application enhances energy efficiency by learning adaptive handover decisions. Once selected, each application independently adjusts its parameters using its own learning policy. More details can be found in [4].
Once an action is taken, the system evaluates its effectiveness by assessing whether the applied changes contribute to the desired network improvements. This evaluation is based on a reward system that provides feedback on whether the selected applications led to better performance. Over time, the system collects a large number of these decision-making instances to form a dataset that contains detailed records of past actions, network conditions, and resulting outcomes. The dataset is obtained via a custom-built simulation environment that mimics real-world network scenarios. Over time, this dataset can be extended and updated with real-world traces and trajectories to enhance its realism and applicability.
The collected data is structured into three distinct datasets. The first dataset is designed for fine-tuning an LLM and consists of query-response pairs and intent-based prompts to enable the model to accurately interpret operator-defined intents and provide relevant network insights. The second dataset comprises time series data, which includes key network parameters such as traffic load, power consumption, and packet loss percentage. This dataset is crucial for predictive intent validation, allowing the system to anticipate future network conditions and assess whether an intent can be successfully executed without degrading performance. The third dataset is a trajectory dataset used in the final stage, where control algorithms learn to optimize network application selection and orchestration. This dataset captures past decisions, network states, and corresponding outcomes to form the basis for training decision models that can make informed, goal-driven choices to achieve operator-defined objectives efficiently.
In this case study, we employ QLoRA to fine-tune an LLM (Llama 3.2 by Meta) to minimize computational resource requirements. Instead of traditional fine-tuning approaches that require high memory and processing power, QLoRA significantly reduces memory consumption without sacrificing performance. This fine-tuned model is further integrated with a RAG module  [^1] The RAG architecture is hybrid: it combines retrieval-based document fetching with the generative capabilities of the fine-tuned LLM., allowing it to access up-to-date network information and respond to dynamic queries with greater accuracy.
In the intent validation stage, the system evaluates whether the operator’s intent can be safely and effectively executed under current and predicted network conditions. This is done using a data-driven and predictive framework. First, the system continuously monitors key performance indicators including traffic load, packet loss percentage, and power consumption. These metrics are forecasted into future time slots using a transformer-based time series model that captures trends and relationships in temporal data. Once an intent is submitted, such as a request to improve throughput or reduce power usage, the system validates it against predicted network conditions. A multi-level check is performed where each intent is compared to predefined thresholds for the three metrics. For example, if a high throughput increase is requested during a period of already high traffic, the system may flag the intent as infeasible. A lookup table is used to cross-reference predicted values and QoS outcomes from past executions to enable fast and intelligent validation. This process helps ensure that only those intents are executed which are likely to maintain or improve overall network performance.
Finally, the last part of this case study includes the Mamba architecture for intent execution via network application orchestration. In particular, we use decision Mamba proposed in [13]. We propose Hierarchical Decision Mamba with Goal Awareness (HDMGA), a hybrid hierarchical decision-making framework that integrates decision Mamba as the high-level decision mechanism and a Decision Transformer (DT) as the low-level control transformer.
At a high level, decision Mamba serves as the goal-aware memory mechanism, responsible for retaining and retrieving critical past actions that have significantly contributed to achieving network optimization objectives. Unlike conventional self-attention mechanisms, which perform exhaustive searches over past trajectories, Mamba employs a Selective-State Space Modeling (SSSM) with a learnable dynamic memory retention mechanism. This allows the system to selectively remember only the most relevant past actions while discarding irrelevant ones. As a result, computational overhead gets reduced and decision efficiency improves.
At the low level, the decision transformer [14] functions as the control transformer, refining real-time actions based on the goal and past knowledge retrieved by the decision Mamba. Once Mamba identifies the most impactful past action, it is passed to the control transformer, which uses a sequence modeling approach to predict the optimal action for the current state. By conditioning decisions on goal awareness rather than predefined reward sequences, the control transformer ensures that each selected action aligns with the operator’s high-level intent.
In this case study, when the operator provides an intent such as “Increase throughput by 10%”, the system sets this increased demand for throughput as its goal. It then recalls a past scenario with similar network conditions, such as high traffic load and moderate packet loss, where a specific combination of applications improved throughput. If the current network state does not support the intent, the system activates a feedback loop that either prompts the operator to revise the intent or uses the LLM to suggest feasible alternatives. The top-level GenAI block (decision Mamba) reviews the goal and the recent network context, retrieves the successful past action, and uses it to guide the present decision.
The control transformer described before has the current state of the network and the past action suggested by the top-level decision Mamba block. It uses this context to decide what to do next. It may decide to take the action of enabling a traffic steering application along with a power control application. This decision is tailored to current needs while being inspired by past success, aiming to fulfill the operator’s intent efficiently.
To evaluate the effectiveness of HDMGA, we compare it against two baseline approaches: Hierarchical Decision Transformer with Goal Awareness (HDTGA) and HRL with intent validation. The HDTGA framework, which is our prior method, utilizes a hierarchical decision transformer architecture. A meta-transformer searches past trajectories to retrieve the most relevant past action, which is then used to guide the control transformer in predicting future actions. The HRL with intent validation baseline follows a conventional HRL approach with a dedicated intent validation mechanism [4].
The simulation setup in this case study consists of a macro cell surrounded by densely deployed small cells in a multi-RAT environment, serving $60$ users. The 5G New Radio (NR) operates at $3.5$ GHz (mid-band) and $30$ GHz (high-band) with bandwidths of $50$ MHz and $100$ MHz, respectively. The max transmission power of the 5G NR BS is $43$ dBm. On the other hand, Long-Term Evolution (LTE) operates at $800$ MHz with a $40$ MHz bandwidth and $38$ dBm maximum transmission power. Traffic types include video, gaming, voice, and vehicle-to-base station data traffic (Ultra-Reliable Low-Latency Communication (URLLC) use case), with packet inter-arrival times of $12.5$ ms, $40$ ms, $20$ ms, and $0.5$ ms, respectively, following Pareto, Uniform, and Poisson distributions.
In Fig. 4, we present spider plots of goal deviations for the proposed HDMGA and baseline methods. Delay-related goals (Fig. 4) include 2, 5, 7, and 11 ms. HDMGA consistently shows the lowest deviation, particularly excelling under stricter constraints. As delay targets relax (e.g., from 7 ms to 11 ms), all methods perform better due to increased flexibility, but HDMGA maintains superior accuracy. Similar trends are seen for energy-efficiency and throughput goals. Compared to HDTGA, HDMGA avoids attention overhead, reducing redundant computations. Additionally, Mamba’s linear-time processing enables efficient handling of long sequences, boosting throughput and making HDMGA more suitable for latency-sensitive applications.
Figure 4: Assessing the GenAI-based proposed three-step methodology in terms of delay goal deviations.
Figure 5: Performance comparison in terms of action inference time.
Fig. 5 presents a plot showing that the action inference time of the proposed method is lower than that of the baselines. The actions illustrated are not sequential decisions taken under identical network conditions. Instead, each action corresponds to an independent inference made under distinct network states at specific time points during the simulation. Since network conditions (e.g., traffic load, power consumption, packet loss) and operator intents vary over time, different methods naturally select varying combinations of applications based on context. This variability reflects each method’s adaptability to dynamic RAN environments. The figure demonstrates that HDMGA achieves faster decision-making, which is essential for near-real-time network automation. This efficiency is enabled by its memory-guided design, where decision Mamba selectively retains and retrieves only the most relevant past actions. This eliminates the need for exhaustive searches over large trajectory buffers. All the methods in the figure maintain inference times within the near-real-time latency bounds defined by O-RAN specifications (10 ms to 1 s) [15].
A major challenge in systems with many network applications is the large and complex action space, which increases training time and hampers generalization. Our prior work addressed this using an attention-based HRL framework to filter irrelevant actions [4]. In this work, the proposed HDMGA further mitigates the issue through supervised learning approach on an offline-collected dataset, eliminating the need for online training during deployment.
V Conclusions
This paper introduces a hierarchical GenAI-driven framework for IDN management in 6G environments by integrating GenAI across intent processing, validation, and execution. By fine-tuning an LLM with QLoRA for efficient intent processing, using transformer-based predictors for validation, and deploying the HDMGA for execution, the proposed approach enhances network automation. The case study demonstrated HDMGA’s superior performance in achieving network objectives like delay reduction, increased energy efficiency, and throughput while reducing computational overhead. In our future works, we wish to integrate RLHF for refining intent validation and execution processes to enable real-time learning and adaptation based on operator feedback and network conditions.
Acknowledgment
This work has been supported by MITACS and Ericsson Canada, and NSERC Canada Research Chairs Program.